A solution is made by diluting 25mL of 0.34 mol/L calcium nitrate, Ca(NO3)2 (aq) solution to 100 mL. Calculate the following concentrations for the solution:
a. the concentration of calcium nitrate
b. the concentration of nitrate ions
1
Expert's answer
2016-01-06T05:36:26-0500
Solution
n(Ca(NO3)2) = V*c = 0.025 * 0.34 = 0.0085 (mol)
c`(Ca(NO3)2) = n/V` = 0.0085 / 0.1 = 0.085 (mol/L)
c`(NO3-) = 2*c(Ca(NO3)2) = 0.17 (mol/L)
Answer
a. c`(Ca(NO3)2) = 0.085 mol/L
b. c`(NO3-) = 0.17 mol/L